# Region-of-Interest-Guided Deep Joint Source-Channel Coding for Image Transmission

### v1

@@ Abstract
Deep joint source-channel coding (deepJSCC) and semantic communication have shown promising improvements in communication performance over wireless networks. However, current approaches primarily focus on enhancing average performance metrics, such as overall image reconstruction quality or task accuracy, which may not fully align with users’ actual experience—often driven by the quality of specific regions of interest (ROI). Motivated by this, we propose ROI-guided joint source-channel coding (ROI-JSCC), a novel deepJSCC framework that prioritizes high-quality transmission of ROI. The ROI-JSCC consists of four key components: (1) ROI embedding and feature map extraction, (2) ROI-guided split processing, (3) ROI-based loss function design, and (4) ROI-adaptive bandwidth allocation. Together, these components enable ROI-JSCC to selectively improve the reconstruction quality of varying ROI while preserving overall image quality without increasing computational burden. Experimental results under diverse communication environments demonstrate that ROI-JSCC significantly improves ROI reconstruction quality while maintaining competitive average image quality compared to recent state-of-the-art methods. All codes are available at https://github.com/hansung-choi/ROI-JSCC.
Index Terms:
joint source-channel coding, region-of-interest, image transmission, quality of experience
@@ I Introduction
With the arrival of the 6G era, a wide range of applications is anticipated, including autonomous driving, smart factories, digital twins, and extended reality technologies such as virtual and augmented reality (VR/AR) [1]. These applications place immense demands on wireless networks, particularly in terms of reliability and latency. However, existing digital communication systems—refined over decades to ensure bit-level accuracy—struggle to meet these requirements. Further improvements to conventional systems have become increasingly difficult, posing significant challenges in supporting the demanding transmission needs of next-generation applications.
@@
To address these challenges, deep learning-based joint source-channel coding (deepJSCC) methods have been proposed for various data modalities, including images [2, 3], point cloud [4], text [5], and speech [6]. These approaches leverage the capability of deep learning models to jointly learn the characteristics of source data and communication channels, enabling the mapping of source data directly into robust and informative channel input symbols. In contrast to traditional communication systems that often fail under severe channel conditions (e.g., high noise levels), deepJSCC methods are able to reliably transmit the meaning of the data. Furthermore, its integration into downstream tasks has led to the development of task-oriented semantic communication (SemCom) systems. By transmitting only task-relevant information, SemCom significantly reduces communication overhead while maintaining high downstream task performance in areas such as image segmentation [7], super-resolution [8], and image classification [9].
Although deepJSCC and SemCom methods have shown great success in preserving the average meaning of data and improving overall task performance, few approaches explicitly consider the ultimate goal of communication in emerging applications. For instance, in autonomous driving, maximizing average semantic segmentation accuracy is not the ultimate objective. Instead, prioritizing performance in important regions, such as pedestrians nearby, even at the cost of reduced segmentation accuracy for less relevant regions like the sky or distant static objects, may contribute more to safe and efficient driving. Similarly, in extended reality, reducing the average reconstruction error between the original and reconstructed images may have a limited impact on the user’s experience. The human visual system is highly sensitive to details in the fovea region, the area near the gaze point, while its sensitivity drops sharply in peripheral regions, the area distant from the gaze point [10]. Therefore, minimizing distortion in the user’s fovea region, even at the expense of decreased quality in peripheral can enhance perceived visual quality. Nevertheless, current deepJSCC and SemCom frameworks lack appropriate coding mechanisms that account for these functional priorities for human experiences.
To address the above issue, we propose a region-of-interest-guided deep joint source-channel coding (ROI-JSCC) for user experience-oriented image transmission. Through the following four mechanisms, the ROI-JSCC effectively preserves high quality in regions of interest (ROI), which dominates user experience: (1) ROI embedding and feature map extraction: This mechanism integrates ROI information into intermediate deepJSCC features, allowing the model to focus on preserving ROI-specific content. (2) ROI-guided split processing: Important ROI features are refined using intensive self-attention [3], while less important features are processed with lightweight spatial attention [11]. This enhances the processing capability for ROI features without incurring significant computational overhead. (3) ROI-based loss function design: Leveraging the mechanism in (1), this loss function guides ROI-JSCC to learn which features are related to ROI and how to encode them to preserve their quality in the transmitted image. (4) ROI-adaptive bandwidth allocation: This mechanism significantly increases bandwidth allocation for ROI features while slightly reducing it for others, thereby enhancing ROI quality with minimal degradation elsewhere.
@@
These mechanisms enable ROI-JSCC to improve ROI-specific image quality across varying ROI configurations without requiring fine-tuning. Moreover, it efficiently utilizes computational resources. Experimental results demonstrate that ROI-JSCC significantly enhances ROI quality while maintaining competitive computational efficiency compared to existing deepJSCC approaches [12, 3, 13, 14].
@@ II ROI-JSCC Communication System
We assume that an image is partitioned into $n_{h}\times n_{w}$ patches. The patches can then be classified into three categories based on their impact on user experience: region of interest (ROI), region of periphery (ROP), and region of non-interest (RONI), as depicted in Figure 4, the ROI map. In this map, ROPs denote patches adjacent to the ROI, while RONIs represent all remaining regions. The importance of a patch decreases as its distance from the ROI increases. We further assume that there is only one ROI and the position of ROI $\gamma:=(h_{\gamma},w_{\gamma})$ in the image is shared between transmitter and receiver, where $h_{\gamma}$ and $w_{\gamma}$ denote the vertical and horizontal coordinates of the ROI in the ROI map, respectively. Our objective is to transmit images in a manner that preserves high quality in ROI, which dominates user experience, while minimizing degradation in overall image quality. To this end, we propose a ROI-JSCC communication system illustrated in Figure 1.
@@
In this system, the encoder $f_{\theta}$ extracts a semantic feature matrix $\mathbf{Z}:=[\mathbf{z}_{1},\ldots,\mathbf{z}_{B}]^{T}\in\mathbb{C}^{B\times C_{m}}$ from a source image $\mathbf{x}\in\mathbb{R}^{H\times W\times 3}$ using the ROI position $\gamma$, where $H$, $W$, and $3$ denote the height, width, and RGB channels of $\mathbf{x}$, respectively. Among the semantic features, some contain significantly more information about the ROI; we refer to these as ROI feature vectors, denoted by $\mathbf{z}_{{\textrm{ROI}}}$. To preserve the information in $\mathbf{z}_{{\textrm{ROI}}}$, the transmitter allocates greater bandwidth to $\mathbf{z}_{{\textrm{ROI}}}$ and less bandwidth to the remaining features by transforming each feature vector dimension, under a total bandwidth constraint. The feature vectors are then aggregated into $\mathbf{\tilde{z}}\in\mathbb{C}^{k}$, where $k$ denotes the total channel bandwidth usage. The channel usage per pixel (CPP) is defined as the ratio of the channel bandwidth to the total number of RGB pixels in $\mathbf{x}$, i.e., ${\textrm{CPP}}:=\frac{k}{3HW}$. Next, the transmitter rescales $\mathbf{\tilde{z}}$ into $\mathbf{\bar{z}}$ to fit the power constraint $P$. Throughout this paper, we assume $P=1$, i.e., $\frac{1}{k}\left\lVert\mathbf{\bar{z}}\right\rVert_{2}^{2}\leq 1$.
@@
After power normalization, $\mathbf{\bar{z}}$ is transmitted over a wireless noisy channel. The receiver then obtains the corrupted feature vector $\mathbf{\hat{z}}_{r}\in\mathbb{C}^{k}$. In this paper, we consider an additive white Gaussian noise (AWGN) channel, modeled as
$\mathbf{\hat{z}}_{r}=\mathbf{\bar{z}}+\mathbf{\epsilon},~{}~{}\mathbf{\epsilon}\sim\mathcal{CN}(0,\sigma^{2}\mathbf{I}_{k})$
where $\mathcal{CN}(0,\sigma^{2}\mathbf{I}_{k})$ denotes a circularly symmetric complex Gaussian distribution with zero mean and covariance matrix $\sigma^{2}\mathbf{I}_{k}$. The quality of the AWGN channel is characterized by the signal-to-noise ratio (SNR), defined as
${\textrm{SNR}}:=10\log_{10}\left(1/\sigma^{2}\right)~{}\text{dB}.$
@@
At the receiver, the received vector $\mathbf{\hat{z}}_{r}\in\mathbb{C}^{k}$ is recovered into $\mathbf{\hat{Z}}^{p}:=[\mathbf{\hat{z}}^{p}_{1},\ldots,\mathbf{\hat{z}}^{p}_{B}]^{T}\in\mathbb{C}^{B\times C_{m}}$ using zero-padding based on the ROI position $\gamma$, as illustrated in Figure 1. The decoder $g_{\phi}$ then reconstructs an image $\mathbf{\hat{x}}$ from $\mathbf{\hat{Z}}^{p}$ and $\gamma$. The quality of the reconstructed image $\mathbf{\hat{x}}$ is evaluated using the peak signal-to-noise ratio (PSNR), which quantifies pixel-wise distortion:
${\textrm{PSNR}}:=10\log_{10}\frac{255^{2}}{{\textrm{MSE}}}~{}\text{dB},$
where MSE denotes the mean squared error between the original and reconstructed images. In this paper, our objective is to improve the PSNR in the ROI (${\textrm{PSNR}}_{{\textrm{ROI}}}$) while maintaining comparable average PSNR (${\textrm{PSNR}}_{{\textrm{Avg}}}$) with respect to previous deepJSCC models. In the following subsections, we introduce the encoder and decoder architectures of ROI-JSCC, the ROI-based loss function, and present the detailed procedure for ROI-adaptive bandwidth allocation.
@@ II-B ROI Block Details
Figure 4: The ROI importance mask generation process.
The detailed structure of the proposed ROI block is illustrated in Figure 3.111For clarity, the refining process is visualized using an image. However, in practice, the ROI block operates on intermediate neural features rather than raw pixel data. First, the ROI importance mask $\gamma_{m}$ is obtained through the generation process depicted in Figure 4. Based on the ROI position $\gamma$, the ROI map is constructed by labeling patches adjacent to the ROI as ROP and all other patches as RONI. Next, the ROI embedding $\gamma_{e}$ is produced by assigning the values $1.0$, $0.5$, and $0.0$ to the ROI, ROP, and RONI patches, respectively. The ROI importance mask $\gamma_{m}$ is then generated by applying the nearest-neighbor method to upsample $\gamma_{e}$ to match the resolution of the ROI block’s input feature $\mathbf{y}_{{\textrm{in}}}$. To effectively incorporate ROI information into $\mathbf{y}_{{\textrm{in}}}$ without degrading its original content, $\gamma_{m}$ is further converted into an importance feature map $\gamma_{f}$, which is then added to $\mathbf{y}_{{\textrm{in}}}$. This integration mechanism enables subsequent layers to identify and emphasize features that are most relevant for improving ${\textrm{PSNR}}_{\textrm{ROI}}$.
@@
As illustrated in Figure 3, $\mathbf{z_{\textrm{ROI}}}$ and the majority of $\mathbf{z_{\textrm{ROP}}}$ are processed using computationally-intensive self-attention [3], whereas $\mathbf{z_{\textrm{RONI}}}$ and the $\mathbf{z_{\textrm{ROP}}}$ vectors adjacent to the three RONI are processed using lightweight spatial attention [11]. Self-attention has proven to be a powerful tool for feature extraction in deepJSCC [3], whereas spatial attention provides satisfactory performance with low computational complexity. Using both mechanisms in a manner guided by ROI, this split processing effectively enhances ${\textrm{PSNR}}_{{\textrm{ROI}}}$ while minimizing degradation in ${\textrm{PSNR}}_{{\textrm{Avg}}}$ without additional computational overhead.
@@ II-C ROI-based Loss Function
Deep neural models have the capacity to learn the relationship between conditional inputs and the loss function. In our setting, the conditional input is $\gamma$, and our goal is to maximize ${\textrm{PSNR}}_{{\textrm{ROI}}}$ while minimizing the degradation in ${\textrm{PSNR}}_{{\textrm{Avg}}}$. To train ROI-JSCC for improving ${\textrm{PSNR}}_{{\textrm{ROI}}}$ across varying $\gamma$, we design the ROI-based loss function $\mathcal{L}$ as follows:
$\mathcal{L}(\mathbf{x},\mathbf{\hat{x}}):=d(\mathbf{x},\mathbf{\hat{x}})+\alpha d(\mathbf{x_{{\textrm{ROI}}}},\mathbf{\hat{x}_{{\textrm{ROI}}}})+\beta d(\mathbf{x_{{\textrm{ROP}}}},\mathbf{\hat{x}_{{\textrm{ROP}}}}),$
where $10<\beta<\alpha<$ 1 are hyperparameters and $d(\cdot)$ denotes the mean squared error (MSE). By using this loss function, ROI-JSCC learns to prioritize features that contribute to improving ${\textrm{PSNR}}_{{\textrm{ROI}}}$ with the aid of $\gamma$, while preserving ${\textrm{PSNR}}_{{\textrm{Avg}}}$. Consequently, it effectively maintains high quality in the ROI during image transmission under varying $\gamma$ without fine tuning.

### v3

@@ Abstract
Deep joint source-channel coding (deepJSCC) methods have shown promising improvements in communication performance over wireless networks. However, existing approaches primarily focus on enhancing overall image reconstruction quality, which may not fully align with user experiences, often driven by the quality of regions of interest (ROI). Motivated by this, we propose ROI-guided joint source-channel coding (ROI-JSCC), a novel deepJSCC framework that prioritizes high-quality transmission of ROI. The ROI-JSCC consists of four key components: (1) Image ROI embedding, (2) ROI-guided split processing, (3) ROI-based loss function design, and (4) ROI-adaptive bandwidth allocation. Together, these components enable ROI-JSCC to selectively enhance the ROI reconstruction quality at varying ROI positions while maintaining overall image quality with minimal computational overhead. Experimental results under diverse communication environments demonstrate that ROI-JSCC significantly improves ROI reconstruction quality while maintaining competitive average image quality compared to recent state-of-the-art methods.
@@ I Introduction
With the arrival of the 6G era, a wide range of applications is anticipated, including autonomous driving, smart factories, digital twins, and extended reality technologies such as virtual and augmented reality (VR/AR) [1]. These applications place immense demands on wireless networks, particularly in terms of reliability and latency. However, conventional digital communication systems—refined over decades to ensure bit-level accuracy—struggle to meet these requirements. Further improvements to conventional systems have become increasingly difficult, posing significant challenges in supporting the demanding transmission needs of next-generation applications.
@@
To address these challenges, deep learning-based joint source-channel coding (deepJSCC) [2, 3] and semantic communication (SemCom) [4] systems have been proposed. DeepJSCC models jointly learn the characteristics of source data and communication channels, enabling the mapping of source data directly into robust and informative channel input symbols. By leveraging the capability of deep learning models, these models have shown high performance in preserving data quality. In contrast, task-oriented SemCom focuses on transmitting task-relevant information, thereby reducing communication overhead while sustaining task performance in applications such as multi-task learning [4]. More recently, region-of-interest-guided semantic communication (ROI-SemCom) methods are proposed to enhance task performance further [5, 6]. These methods identify task-determined ROI areas and allocate more bandwidth to them while assigning less to other areas, thus improving task performance.
However, these methods often fail to account for an essential requirement in real-world applications, namely, the need to reconstruct both the important region and the whole image with sufficient accuracy. For instance, current autonomous driving systems can avoid sudden obstacles using collision prediction, but emergency braking of a leading vehicle often causes rear-end collisions due to insufficient reaction time. Since such accidents dominate autonomous vehicle crashes [7], sharing the scenes where the obstacles are presented more clearly is essential for timely response of following vehicles. In this case, the ROI area containing the suddenly appearing object predicted as high collision probability should be delivered with high accuracy while preserving overall image quality to better support the following car’s decision-making. Although the aforementioned current deepJSCC, SemCom, and ROI-SemCom have achieved remarkable success in overall performance, they fail to meet this goal since they treat the suddenly appearing obstacle and a less relevant distant object in the same way. Moreover, such approaches require retraining or fine-tuning when the ROI position varies.
To address the above issue, we propose a region-of-interest-guided deepJSCC (ROI-JSCC). Our ROI-JSCC focuses on the ROI of images themselves and adaptively enhances the transmission quality of any given image ROIs while preserving average image quality with the following four mechanisms.
(1) Image ROI embedding: This mechanism integrates ROI information into intermediate deepJSCC features, enabling the model to enhance ROI areas. Unlike previous ROI-SemCom embedding methods [5, 6], which rely on task-determined ROIs and cannot handle ROI areas that differ from the task definition, our method can dynamically accommodate varying image ROIs without retraining or fine-tuning.
(2) ROI-guided split processing: Important ROI features are refined using intensive self-attention [3], while less important features are processed with lightweight spatial attention [8]. Unlike previous deepJSCC that applies intensive self-attention for all features without considering ROI importance [3], ours is the first computational mechanism that applies self-attention to the only ROI related features. This efficiently enhances the processing capability for ROI features with low computational burden.
(3) ROI-based loss function design: Leveraging the mechanism in (1), a new loss function guides ROI-JSCC to learn which features are related to ROI and how to encode them to preserve their quality in the transmitted image. Unlike previous ROI-SemCom loss functions, which are designed to enhance task-determined ROIs, our loss function enables the model to enhance dynamically varying image ROIs without retraining.
(4) ROI-adaptive bandwidth allocation: This mechanism significantly increases bandwidth allocation for ROI features while slightly reducing it for others, thereby enhancing ROI quality with minimal degradation elsewhere. The previous ROI-SemCom methods [5, 6] need additional transmission to inform the bandwidth allocation result of every feature pixel. In contrast, our method only needs to transmit a few bits of the ROI’s coordinate location, which significantly reduces the extra transmission burden.
@@
These mechanisms enable ROI-JSCC to improve ROI-specific image quality across varying ROI configurations without requiring retraining or fine-tuning. Moreover, it efficiently utilizes computational resources by allocating more resources to the ROI area.
@@ II ROI-JSCC Communication System
We first partition the image into $n_{h}\times n_{w}$ patches, among which our primary focus is a single patch representing the region-of-interest (ROI). We assume that the position of ROI $\gamma:=(h_{\gamma},w_{\gamma})$ in the image is known to the transmitter and receiver, where $h_{\gamma}$ and $w_{\gamma}$ denote the vertical and horizontal indices in the ROI map, see Figure 4.111In the driving example in Section I, the transmitter identifies the obstacle using an algorithm, e.g., segmentation, and determines the ROI’s location with obstacle’s collision probability. This ROI position needs to be shared with the receiver; however, since it requires only $\log_{2}(n_{h}n_{w})$ bits, negligible relative to the total bandwidth usage, we ignore this communication overhead. For example, when $n_{h}=n_{w}=4$, an additional $4$ bits transmission is required. Thus, we can assume that the ROI position is available to both the transmitter and the receiver. More detailed explanations are provided in Appendix A. Adjacent patches are categorized as region of periphery (ROP) or region of non-interest (RONI) according to their relative location with respect to the ROI, see Figure 4 for example. Our objective is to transmit images in a manner that preserves high quality in ROI, which dominates user experience, while minimizing degradation in overall image quality at the same time. To this end, we propose a ROI-JSCC communication system illustrated in Figure 1.
@@
In this system, the encoder $f_{\theta}$ extracts a feature matrix $\mathbf{Z}:=[\mathbf{z}_{1},\ldots,\mathbf{z}_{B}]^{T}\in\mathbb{C}^{B\times C_{m}}$ from a source image $\mathbf{x}\in\mathbb{R}^{H\times W\times 3}$ using the ROI position $\gamma$, where $H$, $W$, and $3$ denote the height, width, and RGB channels of $\mathbf{x}$, respectively. Among the features, some contain significantly more information about the ROI; we refer to these as ROI feature vectors, denoted by $\mathbf{z}_{{\textrm{ROI}}}$. To preserve the information in $\mathbf{z}_{{\textrm{ROI}}}$, the transmitter allocates greater bandwidth to $\mathbf{z}_{{\textrm{ROI}}}$ and less bandwidth to the remaining features by transforming each feature vector dimension, under a total bandwidth constraint. The feature vectors are then aggregated into $\mathbf{\tilde{z}}\in\mathbb{C}^{k}$, where $k$ denotes the total channel bandwidth usage. The channel usage per pixel (CPP) is defined as the ratio of the channel bandwidth to the total number of RGB pixels in $\mathbf{x}$, i.e., ${\textrm{CPP}}:=\frac{k}{3HW}$. Next, the transmitter rescales $\mathbf{\tilde{z}}$ into $\mathbf{\bar{z}}$ to fit the power constraint $P$. Throughout this paper, we assume $P=1$, i.e., $\frac{1}{k}\left\lVert\mathbf{\bar{z}}\right\rVert_{2}^{2}\leq 1$.
@@
After power normalization, $\mathbf{\bar{z}}$ is transmitted over a wireless noisy channel. The receiver then obtains the corrupted feature vector $\mathbf{\hat{z}}_{r}\in\mathbb{C}^{k}$. In this paper, we consider an additive white Gaussian noise (AWGN) and a fast Rayleigh fading channel, modeled as $\hat{z}_{r,i}=h_{i}\bar{z}_{i}+\epsilon_{i},~~\mathbf{\epsilon_{i}}\sim\mathcal{CN}(0,\sigma^{2})$ where $\mathcal{CN}(0,\sigma^{2})$ denotes a complex Gaussian distribution and subscript $i$ denotes $i$-th symbol of the corresponding vector. Here, $h_{i}=1$ for the AWGN channel and $h_{i}\sim\mathcal{CN}(0,1)$ for the fast Rayleigh fading channel. The quality of the communication channels is characterized by the signal-to-noise ratio (SNR), defined as ${\textrm{SNR}}:=10\log_{10}\left(1/\sigma^{2}\right)~\text{dB}.$
@@
At the receiver, the received vector $\mathbf{\hat{z}}_{r}\in\mathbb{C}^{k}$ is recovered into $\mathbf{\hat{Z}}^{p}:=[\mathbf{\hat{z}}^{p}_{1},\ldots,\mathbf{\hat{z}}^{p}_{B}]^{T}\in\mathbb{C}^{B\times C_{m}}$ using zero-padding based on the ROI position $\gamma$, as illustrated in Figure 1. The decoder $g_{\phi}$ then reconstructs an image $\mathbf{\hat{x}}$ from $\mathbf{\hat{Z}}^{p}$ and $\gamma$. The quality of the reconstructed image $\mathbf{\hat{x}}$ is evaluated using the peak signal-to-noise ratio (PSNR) and the structural similarity index measure (SSIM). The PSNR quantifies pixel-wise distortion: ${\textrm{PSNR}}:=10\log_{10}\frac{255^{2}}{{\textrm{MSE}}(\mathbf{x},\mathbf{\hat{x}})}~\text{dB},$ and the SSIM is a perceptual metric: ${\textrm{SSIM}}(\mathbf{x},\hat{\mathbf{x}}):=l(\mathbf{x},\hat{\mathbf{x}})^{p}c(\mathbf{x},\hat{\mathbf{x}})^{q}s(\mathbf{x},\hat{\mathbf{x}})^{r},$ where $l(\cdot)$, $c(\cdot)$, and $s(\cdot)$ measure brightness, contrast, and structural similarities and $p,q,r$ weight contributions of each measure. In this paper, our objective is to improve the performance in the ROI (${\textrm{PSNR}}_{{\textrm{ROI}}}$, ${\textrm{SSIM}}_{{\textrm{ROI}}}$) while maintaining comparable average performance (${\textrm{PSNR}}_{{\textrm{Avg}}}$, ${\textrm{SSIM}}_{{\textrm{Avg}}}$) with respect to previous deepJSCC models. In the following subsections, we introduce the encoder and decoder architectures of ROI-JSCC, the ROI-based loss function, and present the detailed procedure for ROI-adaptive bandwidth allocation.
@@ II-B ROI Block Details
Figure 4: The ROI importance mask generation process for ROI $\gamma=(2,2)$.
The detailed structure of the proposed ROI block is illustrated in Figure 3.222To illustrate, the refinement process is visualized using an image, but the ROI block in fact operates on intermediate neural features rather than raw pixel data. First, the ROI importance mask $\gamma_{m}$ is obtained through the generation process depicted in Figure 4. Based on the ROI position $\gamma$, the ROI map is constructed by labeling patches adjacent to the ROI as ROP and all other patches as RONI. Next, the ROI embedding $\gamma_{e}$ is produced by assigning the values $1.0$, $0.5$, and $0.0$ to the ROI, ROP, and RONI patches, respectively. The ROI importance mask $\gamma_{m}$ is then generated by applying the nearest-neighbor method to upsample $\gamma_{e}$ to match the resolution of the ROI block’s input feature $\mathbf{y}_{{\textrm{in}}}$. To effectively incorporate ROI information into $\mathbf{y}_{{\textrm{in}}}$ without degrading its original content, $\gamma_{m}$ is further converted into an importance feature map $\gamma_{f}$, which is then added to $\mathbf{y}_{{\textrm{in}}}$. This integration mechanism enables subsequent layers to identify and emphasize features that are most relevant for improving ${\textrm{PSNR}}_{\textrm{ROI}}$.
@@
As illustrated in Figure 3, $\mathbf{z_{\textrm{ROI}}}$ and the majority of $\mathbf{z_{\textrm{ROP}}}$ are processed using computationally-intensive self-attention [3], whereas $\mathbf{z_{\textrm{RONI}}}$ and the $\mathbf{z_{\textrm{ROP}}}$ vectors adjacent to the three RONI are processed using lightweight spatial attention [8]. Self-attention has proven to be a powerful tool for feature extraction in deepJSCC [3], whereas spatial attention provides satisfactory performance with low computational complexity. By leveraging both mechanisms under ROI guidance, the proposed split processing not only enhances ${\textrm{PSNR}}_{{\textrm{ROI}}}$ but also minimizes degradation in ${\textrm{PSNR}}_{{\textrm{Avg}}}$, which is the first efficient mechanism that allocates different computational resources to ROI and non-ROI regions.
@@ II-C ROI-based Loss Function
Deep neural models have the capacity to learn the relationship between conditional inputs and the loss function. In our setting, the conditional input is $\gamma$, and our goal is to maximize the image quality of the ROI region while minimizing the average performance degradation. To train ROI-JSCC for improving ${\textrm{PSNR}}_{{\textrm{ROI}}}$ across varying $\gamma$, we design the ROI-based loss function $\mathcal{L}$ as follows: $\mathcal{L}(\mathbf{x},\mathbf{\hat{x}}):=d(\mathbf{x},\mathbf{\hat{x}})+\alpha d(\mathbf{x_{{\textrm{ROI}}}},\mathbf{\hat{x}_{{\textrm{ROI}}}})+\beta d(\mathbf{x_{{\textrm{ROP}}}},\mathbf{\hat{x}_{{\textrm{ROP}}}}),$ where $0<\beta<\alpha<1$ are hyperparameters and $d(\cdot)$ is the mean squared error (MSE) for the PSNR metric and $1-{\textrm{SSIM}}$ for the SSIM metric. By using this loss function, ROI-JSCC learns to prioritize features that contribute to improving the ROI region quality with the aid of $\gamma$, while preserving average quality. Consequently, it effectively maintains high quality in the ROI during image transmission under varying $\gamma$ without fine-tuning.